# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multimodal Large Language Models

### v5

@@ Abstract
1 Introduction
@@
Deep generative models, such as Variational Autoencoders (VAEs) Kingma and Welling (2013) and diffusion models Rombach et al. (2022), have become a state-of-the-art approach in various generation tasks Ho et al. (2022); Yang et al. (2023). These methods effectively leverage latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Inductive biases are often enforced over latent variables in deep generative models. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables Ding et al. (2020). Moreover, sometimes latent variables can be grouped, leading to combination bias Klys et al. (2018). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest Wang et al. (2024), has given rise to conditional bias. Incorporating inductive biases aligned with the actual facts can reduce the hallucination in explaining the latent variables in deep generative models Wu et al. (2024).
@@
Figure 1: Illustration of how a pretrained generative model, guided by inductive bias formulas, automatically decodes and interprets meaningful latent dimension semantics.
@@
The field of explainable artificial intelligence (XAI) has extensively investigated the interpretation of machine learning models Adadi and Berrada (2018); Zhu et al. (2021). However, interpreting latent variables in deep generative models remains underexplored. Machine learning model explanation, a.k.a., post-hoc explanation, can be categorized into global and local explanations Gao et al. (2024). Global explanations focus on elucidating the entire model, while local explanations target the reasoning behind specific predictions. Global explanations are more challenging, with existing work mostly emphasizing attributions to identify which features are most important for model decision-making Saleem et al. (2022). The missing piece is understanding the meaning of features when they are unknown, which is very common in deep generative models. More recently, one category of global explanation methods, called concept-based explanations, aims to generate more human-understandable concepts as explanations Poeta et al. (2023). However, current concept-based methods often rely on human heuristics or predefined concept and feature space, limiting the expressiveness of the explanations and falling short of achieving truly automatic explanation generation Koh et al. (2020); Bai et al. (2022).
@@
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models (addressing Challenge 1), we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (addressing Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables (addressing Challenge 3), we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent representations are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
2 Related Work
@@
Deep Generative Models. Deep generative models are essential for modeling complex data distributions. Variational Autoencoders (VAEs) are prominent in this area, introduced by Kingma and Welling Kingma and Welling (2013). VAEs encode input data into a latent space and decode it back, optimizing a balance between reconstruction error and the Kullback-Leibler divergence Rezende et al. (2014). They have diverse applications, including image generation Yan et al. (2016), and anomaly detection An and Cho (2015).
@@
Diffusion models, proposed by Ho et al. Ho et al. (2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples Rombach et al. (2022). Advances have made them applicable to text-to-image synthesis Nichol and Dhariwal (2021), and audio generation Kong et al. (2020)
@@
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments Chen et al. (2016). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images Brock et al. (2016); Zhu et al. (2016). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations Shen et al. (2020); Esser et al. (2020). Some concept-based models control latent variables in generative models using category concepts to generate data Tran et al. (2022); Bouchacourt et al. (2018). However, these approaches cannot automatically generate free-form textual explanations.
@@
More recently, Multimodal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information Yin et al. (2023); Bai et al. (2024). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities OpenAI (2024), and Gemini that is a family of highly capable multimodal models Team et al. (2024). We envision leveraging MLLMs to automatically generate explanations for latent variables and incorporate their inductive biases to reduce hallucination. Our work focuses on: (1) how to decompose the inductive bias formulas to automate the manipulation of latent variables, (2) how to develop prompts that are aligned with the underlying inductive biases of generative models and can be easily understood by MLLMs, and (3) how to evaluate the quality of the generated explanations for latent variables.
@@
Figure 2: The overview of our proposed framework LatentExplainer (a) Inductive-bias-guided Data Manipulation generates image sequences by manipulating latent variables with predefined biases; (b) Automatic Prompt Generation with symbol-to-word mapping uses these images and formulas to create prompts for an MLLM to produce explanations; (c) Uncertainty-aware Explanation Generation evaluates multiple responses from the MLLM, selecting the most consistent explanation with a certainty score.
3 Preliminaries and Problem Formulation
@@
Deep Generative Models. These are a class of models that learn a mapping between observations $\mathbf{x}$ and key underlying factors $\mathbf{z}$. These models are widely used in deep generative models such as VAEs and latent diffusion models. VAEs, for instance, introduce a probabilistic approach to encoding data by maximizing the evidence lower bound (ELBO) Kingma and Welling (2013):
@@
Latent diffusion models further refine the generation process by iteratively refining noise into structured data Rombach et al. (2022). These models effectively capture the underlying structure of data in a low-dimensional semantic space. Latent variable manipulation of diffusion models aims at transversing the latent representation $\mathbf{z}$ along the semantic latent direction $\mathbf{v}$. The perturbed vector $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is a hyper-parameter controlling the strength and $\mathcal{G}$ is a diffusion decoder Park et al. (2023). An image sequence can then be generated by $\mathcal{G}(\mathbf{\tilde{z}})$. The perturbations in the semantic latent direction lead to semantic changes in the generated image sequence.
@@
Inductive Bias in Latent Variables. Inductive biases are usually imposed on latent variables to enhance the performance and interpretability of deep generative models. These inductive biases in deep generative models can be categorized into three common types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data Ding et al. (2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures Klys et al. (2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables Wang et al. (2024).
@@
Problem Formulation. We assume a dataset $\mathcal{D}$, where each sample consists of $x$ or $(x,y)$, with $x\in\mathbb{R}^{N}$ and $y=\left\{y_{k}\in\right.$$\mathbb{R}\}_{k=1}^{K}$ as $K$ properties of $x$. The dataset $\mathcal{D}$ is generated by M latent variables $z_{i}$, where $i\in\{1,\ldots,M\}$. $z_{i}$ can be a single latent variable or a group of correlated latent variables. Suppose we are given a generative model with a set of formulas $\mathcal{F}$ with respect to $z_{i}$, where $\mathcal{F}$ represents an inductive bias that the generative model must satisfy. Our goal is to derive a textual sequence that explains the semantic meanings of the latent variable $z_{i}$.
4 Proposed Method
4.1 Overview of LatentExplainer
@@
This paper focuses on the tasks in explaining the semantics of latent variables $\{{z}_{i}\}_{i=1}^{M}$ in deep generative models. To interpret the semantics of latent variables and work around the blackbox nature of deep generative models, we propose to perturb each latent variable and explain the change it imposes on the generated data. To solve these challenges, we propose a novel LatentExplainer scheme. The pseudo-code of this whole scheme can be found in Algorithm 1. When explaining latent variable models, it is crucial to fully leverage and align with the prior knowledge about them. To do this, we design a generic framework that can automatically formulate inductive bias of generative models into textual prompts. Specifically, we have summarized three common inductive biases and designed their symbol-to-word one-on-one prompts $\mathcal{P}$ (Section 4.2). Our scheme can adaptively convert a user-provided inductive bias formulas $\mathcal{F}$ into a corresponding prompt $\mathcal{P}$ to provide more accurate explanations of the latent representations in Figure 2(b) (Section 4.3). We decompose the inductive bias to guide the perturbation of $z_{i}$ and subsequently decode the manipulated latent variables into generated data that are perceptible by humans such as images. Through a series of perturbations on $z_{i}$, a sequence of generated data samples can be obtained to reflect the changes in $z_{i}$ in Figure 2(a) (Section 4.4). Eventually, the explanations are selected through an uncertainty quantification approach to assess whether the latent representations are interpretable and select the most consistent explanations in Figure 2(c) (Section 4.5).
4.2 Inductive-bias-guided Prompt Framework
4.2.1 Generic Framework
@@ 8 $\neq$ change
Table 1: Lookup table for symbol-to-word mapping.
Fixed Ending: What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation”.
4.2.2 From Disentanglement Bias to Prompts
Disentanglement bias refers to the model’s ability to separate independent factors in the data Ding et al. (2020); Wu et al. (2023). The formula representing this bias focuses on ensuring that different latent variables correspond to different independent underlying factors. Independent factors would be invariant with respect to one another Ridgeway (2016). By disentangling these factors, researchers can better understand the underlying structure of the data and improve the model’s performance on tasks such as representation learning.
@@ 4.2.2 From Disentanglement Bias to Prompts
4.2.3 From Combination Bias to Prompts
Combination bias involves understanding how different latent variables interact within groups and remain independent across groups Klys et al. (2018). This bias is significant as it helps in identifying how combinations of factors contribute to the overall data generation process. Recognizing these interactions enables researchers to design models that can generate more complex and realistic data by capturing intricate relationships within the data.
@@ 4.2.3 From Combination Bias to Prompts
4.2.4 From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables Wang et al. (2024). This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@ 4.2.4 From Conditional Bias to Prompts
4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning. A more detailed example is at Appendix D.
As Algorithm 1 shows, our approach starts by extracting mathematical symbols from the given formula $\mathcal{F}$ using the ExtractSymbols function (line 4). This function traverses the formula to identify the mathematical symbols.

### v6

@@ Abstract.
Explainable AI, Multimodal Large Language Models, Deep Generative Models, Latent Representations
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; May; 2025isbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Deep generative models, such as Variational Autoencoders (VAEs) (Kingma and Welling, 2013) and diffusion models (Rombach et al., 2022), have become a state-of-the-art approach in various generation tasks (Ho et al., 2022; Yang et al., 2023). These methods effectively leverage latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Inductive biases are often enforced over latent variables in deep generative models. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables (Ding et al., 2020). Moreover, sometimes latent variables can be grouped, leading to combination bias (Klys et al., 2018). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest (Wang et al., 2024), has given rise to conditional bias. Incorporating inductive biases aligned with the actual facts can reduce the hallucination in explaining the latent variables in deep generative models (Wu et al., 2024).
@@
Figure 1. Illustration of how a pretrained generative model, guided by inductive bias formulas, automatically decodes and interprets meaningful latent dimension semantics.
@@
The field of explainable artificial intelligence (XAI) has extensively investigated the interpretation of machine learning models (Adadi and Berrada, 2018; Zhu et al., 2021). However, interpreting latent variables in deep generative models remains underexplored. Machine learning model explanation, a.k.a., post-hoc explanation, can be categorized into global and local explanations (Gao et al., 2024). Global explanations focus on elucidating the entire model, while local explanations target the reasoning behind specific predictions. Global explanations are more challenging, with existing work mostly emphasizing attributions to identify which features are most important for model decision-making (Saleem et al., 2022). The missing piece is understanding the meaning of features when they are unknown, which is very common in deep generative models. More recently, one category of global explanation methods, called concept-based explanations, aims to generate more human-understandable concepts as explanations (Poeta et al., 2023). However, current concept-based methods often rely on human heuristics or predefined concept and feature space, limiting the expressiveness of the explanations and falling short of achieving truly automatic explanation generation (Koh et al., 2020; Bai et al., 2022).
@@
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models (addressing Challenge 1), we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (addressing Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables (addressing Challenge 3), we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent variables are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
2. Related Work
@@
Deep Generative Models. Deep generative models are essential for modeling complex data distributions. Variational Autoencoders (VAEs) are prominent in this area, introduced by Kingma and Welling (Kingma and Welling, 2013). VAEs encode input data into a latent space and decode it back, optimizing a balance between reconstruction error and the Kullback-Leibler divergence (Rezende et al., 2014). They have diverse applications, including image generation (Yan et al., 2016), and anomaly detection (An and Cho, 2015).
@@
Diffusion models, proposed by Ho et al. (Ho et al., 2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing generative adversarial networks (GANs) in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples (Rombach et al., 2022). Advances have made them applicable to text-to-image synthesis (Nichol and Dhariwal, 2021), and audio generation (Kong et al., 2020)
@@
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments (Chen et al., 2016). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images (Brock et al., 2016; Zhu et al., 2016). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations (Shen et al., 2020; Esser et al., 2020). Some concept-based models control latent variables in generative models using category concepts to generate data (Tran et al., 2022; Bouchacourt et al., 2018). However, these approaches cannot automatically generate free-form textual explanations.
@@
More recently, Multimodal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information (Yin et al., 2023; Bai et al., 2024). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities (OpenAI, 2024), and Gemini that is a family of highly capable multimodal models (Team et al., 2024). We envision leveraging MLLMs to automatically generate explanations for latent variables and incorporate their inductive biases to reduce hallucination. Our work focuses on: (1) how to decompose the inductive bias formulas to automate the manipulation of latent variables, (2) how to develop prompts that are aligned with the underlying inductive biases of generative models and can be easily understood by MLLMs, and (3) how to evaluate the quality of the generated explanations for latent variables.
@@
Figure 2. The overview of our proposed framework LatentExplainer (a) Inductive-bias-guided Data Manipulation generates image sequences by manipulating latent variables with predefined biases; (b) Automatic Prompt Generation with symbol-to-word mapping uses these images and formulas to create prompts for an MLLM to produce explanations; (c) Uncertainty-aware Explanation Generation evaluates multiple responses from the MLLM, selecting the most consistent explanation with a certainty score.
3. Preliminaries and Problem Formulation
@@
Deep Generative Models. These are a class of models that learn a mapping between observations $\mathbf{x}$ and underlying latent factors $\mathbf{z}$. These models are widely used in deep generative models such as VAEs and latent diffusion models. VAEs, for instance, introduce a probabilistic approach to encoding data by maximizing the evidence lower bound (ELBO) (Kingma and Welling, 2013):
@@
Latent diffusion models further refine the generation process by iteratively refining noise into structured data (Rombach et al., 2022). These models effectively capture the underlying structure of data in a low-dimensional semantic space. Latent variable manipulation of diffusion models aims at transversing the latent variables $\mathbf{z}$ along the semantic latent direction $\mathbf{z_{i}}$. The perturbed vector $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{z_{i}})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is a hyper-parameter controlling the strength and $\mathcal{G}$ is a diffusion decoder (Park et al., 2023). An image sequence can then be generated by $\mathcal{G}(\mathbf{\tilde{z}})$. The perturbations in the semantic latent direction lead to semantic changes in the generated image sequence.
@@
Inductive Bias in Latent Variables. Inductive biases are usually imposed on latent variables to enhance the performance and interpretability of deep generative models. These inductive biases in deep generative models can be categorized into three common types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data (Ding et al., 2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures (Klys et al., 2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables (Wang et al., 2024).
@@
Problem Formulation. We assume a dataset $\mathcal{D}$, where each sample consists of $x$ or $(x,y)$, with $x\in\mathbb{R}^{N}$ and $y=\left\{y_{k}\in\right.$$\mathbb{R}\}_{k=1}^{K}$ as $K$ properties of $x$. The dataset $\mathcal{D}$ is generated by M latent variables $z_{i}$, where $i\in\{1,\ldots,M\}$. $z_{i}$ can be an individual latent dimension in $z$ in the VAE models or a latent vector in the diffusion models. Suppose we are given a generative model with a set of formulas $\mathcal{F}$ with respect to $z_{i}$, where $\mathcal{F}$ represents an inductive bias that the generative model must satisfy. Our goal is to derive a textual sequence that explains the semantic meanings of $z_{i}$.
4. Proposed Method
4.1. Overview of LatentExplainer
@@
This paper focuses on the tasks in explaining the semantics of latent variables $\{{z}_{i}\}_{i=1}^{M}$ in deep generative models. To interpret the semantics of latent variables and work around the blackbox nature of deep generative models, we propose to perturb each $z_{i}$ and explain the change it imposes on the generated data. To solve these challenges, we propose a novel LatentExplainer scheme. The pseudo-code of this whole scheme can be found in Algorithm 1. When explaining latent variable models, it is crucial to fully leverage and align with the prior knowledge about them. To do this, we design a generic framework that can automatically formulate inductive bias of generative models into textual prompts. Specifically, we have summarized three common inductive biases and designed their symbol-to-word one-on-one prompts $\mathcal{P}$ (Section 4.2). Our scheme can adaptively convert a user-provided inductive bias formulas $\mathcal{F}$ into a corresponding prompt $\mathcal{P}$ to provide more accurate explanations of the latent variables in Figure 2(b) (Section 4.3). We decompose the inductive bias to guide the perturbation of $z_{i}$ and subsequently decode the manipulated latent variables into generated data that are perceptible by humans such as images. Through a series of perturbations on $z_{i}$, a sequence of generated data samples can be obtained to reflect the changes in $z_{i}$ in Figure 2(a) (Section 4.4). Eventually, the explanations are selected through an uncertainty quantification approach to assess whether the latent variables are interpretable and select the most consistent explanations in Figure 2(c) (Section 4.5).
4.2. Inductive-bias-guided Prompt Framework
4.2.1. Generic Framework
@@ 8 $\neq$ change
Table 1. Lookup table for symbol-to-word mapping.
Fixed Ending: What is the pattern of change? Write in a sentence. If there is no clear pattern, just write ”No clear explanation”.
4.2.2. From Disentanglement Bias to Prompts
Disentanglement bias refers to the model’s ability to separate independent factors in the data (Ding et al., 2020; Wu et al., 2023). The formula representing this bias focuses on ensuring that different latent variables correspond to different independent underlying factors. Independent factors would be invariant with respect to one another (Ridgeway, 2016). By disentangling these factors, researchers can better understand the underlying structure of the data and improve the model’s performance on tasks such as representation learning.
@@ 4.2.2. From Disentanglement Bias to Prompts
4.2.3. From Combination Bias to Prompts
Combination bias involves understanding how different latent variables interact within groups and remain independent across groups (Klys et al., 2018). This bias is significant as it helps in identifying how combinations of factors contribute to the overall data generation process. Recognizing these interactions enables researchers to design models that can generate more complex and realistic data by capturing intricate relationships within the data.
@@ 4.2.3. From Combination Bias to Prompts
4.2.4. From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables (Wang et al., 2024). This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
@@ 4.2.4. From Conditional Bias to Prompts
4.3. Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ within these three bias using in-context learning. A more detailed example is at Appendix A.
As Algorithm 1 shows, our approach starts by extracting mathematical symbols from the formula $\mathcal{F}$ using the ExtractSymbols function (line 4). This function traverses the formula to identify the mathematical symbols.